# Channel Capacity of Near-Field Multiuser Communications

### v1

@@ III-D Extension to Cases of $K>2$
The sum-rate capacity of the MAC for $K>2$ is given by
${\mathsf{C}}_{\mathsf{MAC}}=\log_{2}\det\Big{(}{\mathbf{I}}+{\sigma^{-2}}\sum\nolimits_{k=1}^{K}{P_{k}}{\mathbf{h}}_{k}{\mathbf{h}}_{k}^{\mathsf{H}}\Big{)}.$ (33)
Following steps similar to those in Appendix -A, (33) can be rewritten as $\mathsf{C}_{\mathsf{MAC}}=\log_{2}\det\left(\mathbf{A}\right)$ with
$\mathbf{A}=\left[\begin{matrix}1+\frac{P_{1}}{\sigma^{2}}\|\mathbf{h}_{1}\|^{2}&\cdots&1+\frac{\sqrt{P_{1}P_{K}}}{\sigma^{2}}\mathbf{h}_{1}^{\mathsf{H}}\mathbf{h}_{K}\\ \vdots&\ddots&\vdots\\ 1+\frac{\sqrt{P_{1}P_{K}}}{\sigma^{2}}\mathbf{h}_{K}^{\mathsf{H}}\mathbf{h}_{1}&\cdots&1+\frac{P_{K}}{\sigma^{2}}\|\mathbf{h}_{K}\|^{2}\\ \end{matrix}\right],$ (34)
which can be expressed in terms of a function of the channel gains and CCFs.
The capacity region of the MAC exhibits a polymatroidal structure in $K$-dimensional space, and each of the $K!$ corner points can be attained by employing SIC decoding in a specific message decoding order. The remaining capacity region is realized through time-sharing among these corner points. Specifically, when considering the SIC decoding order $\varpi(K)\rightarrow\varpi(K-1)\rightarrow\ldots\rightarrow\varpi(1)$ with $\{\varpi(k)\}_{k=1}^{K}={\mathcal{K}}$, the achievable rate tuple $({\mathsf{R}}_{1},\ldots,{\mathsf{R}}_{K})$ at each corner point of the capacity region can be calculated as
$\begin{split}&\mathsf{R}_{\varpi\left(k\right)}=\log_{2}\left(1+{P_{\varpi(k)}}{\sigma^{-2}}\|\mathbf{h}_{\varpi(k)}\|^{2}+\right.\\ &\times\left.\frac{P_{1}P_{2}\sigma^{-4}\sum\nolimits_{k^{\prime}<k}{P_{\varpi(k)}P_{\varpi(k^{\prime})}|\mathbf{h}_{\varpi(k)}^{\mathsf{H}}\mathbf{h}_{\varpi(k^{\prime})}|^{2}}}{1+\sigma^{-2}\sum\nolimits_{k^{\prime}<k}{P_{\varpi(k^{\prime})}\|\mathbf{h}_{\varpi(k^{\prime})}\|^{2}}}\right),\end{split}$ (35)
which is also a function of the channel gains and CCFs. Therefore, the results derived in the case of $K=2$ can be trivially extended to the case of $K>2$.
@@ IV-A Sum-Rate Capacity
${\mathsf{R}}_{\varepsilon(k)}=\log_{2}\left(1\!+\!\frac{{\mathbf{h}}_{\varepsilon(k)}^{\mathsf{H}}{\bm{\Sigma}}_{\varepsilon(k)}{\mathbf{h}}_{\varepsilon(k)}}{\sigma_{\varepsilon(k)}^{2}\!+\!\sum_{k^{\prime}<k}{\mathbf{h}}_{\varepsilon(k^{\prime})}^{\mathsf{H}}{\bm{\Sigma}}_{\varepsilon(k^{\prime})}{\mathbf{h}}_{\varepsilon(k^{\prime})}}\right).$ (36)
@@
${\mathsf{C}}_{\mathsf{BC}}=$ $\max_{\{{\bm{\Sigma}}_{k}\}_{k=1}^{2}:{\bm{\Sigma}}_{k}\succeq{\mathbf{0}},\sum_{k=1}^{2}{\mathsf{tr}}({\bm{\Sigma}}_{k})\leq P}\log_{2}\left(1+\frac{{\mathbf{h}}_{1}^{\mathsf{H}}{\bm{\Sigma}}_{1}{\mathbf{h}}_{1}}{\sigma_{1}^{2}}\right)$
$\qquad\qquad\qquad+\log_{2}\left(1+\frac{{\mathbf{h}}_{2}^{\mathsf{H}}{\bm{\Sigma}}_{2}{\mathbf{h}}_{2}}{\sigma_{2}^{2}+{\mathbf{h}}_{2}^{\mathsf{H}}{\bm{\Sigma}}_{1}{\mathbf{h}}_{2}}\right).$ (38)
@@
Problem (IV-A) is a non-convex problem. Thus, numerically finding the maximum is a nontrivial problem. However, in [22], a duality is shown to exist between the uplink and downlink which establishes that the dirty-paper capacity region for the BC is equal to the capacity region of the dual MAC (described in (7)). Therefore, (IV-A) can be rewritten as
$\!{\mathsf{C}}_{\mathsf{BC}}$ $=\max_{0\leq p_{k},\sum_{k=1}^{2}p_{k}=P}\log_{2}\det\left({\mathbf{I}}+\sum\nolimits_{k=1}^{K}\frac{p_{k}}{\sigma_{k}^{2}}{\mathbf{h}}_{k}{\mathbf{h}}_{k}^{\mathsf{H}}\right)$
$\!\!\!=\max_{0\leq p_{k},\sum_{k=1}^{2}p_{k}=P}\!\log_{2}\!\left(1\!+\frac{p_{1}}{\sigma_{1}^{2}}\left\|\mathbf{h}_{1}\right\|^{2}\!+\frac{p_{2}}{\sigma_{2}^{2}}\left\|\mathbf{h}_{2}\right\|^{2}\right.$
$\qquad\qquad\qquad\quad\left.+\frac{p_{1}p_{2}}{\sigma_{1}^{2}\sigma_{2}^{2}}\left\|\mathbf{h}_{1}\right\|^{2}\left\|\mathbf{h}_{2}\right\|^{2}\left(1-\rho\right)\right).$ (39)
Note that problem (IV-A) is a convex problem which can be optimally solved. After obtaining the optimized $\{p_{k}\}_{k=1}^{2}$, we can employ the encoding order as well as the transformation in [22, Equs. (8)–(10)] and [23, Appendix A] to recover the corresponding downlink covariance matrices that achieve the same rates and the sum-rate.
@@
The optimal power allocation in the NF scenario is given by
$\left(p_{1}^{\star},p_{2}^{\star}\right)=\begin{cases}\left(P,0\right)&\kappa_{1}\leq 0\\ \left(0,P\right)&\kappa_{2}\leq 0\\ \left(\frac{\kappa_{1}\sigma_{2}^{2}}{G_{2}\bar{\rho}},\frac{\kappa_{2}\sigma_{1}^{2}}{G_{1}\bar{\rho}}\right)&\mathrm{else}\\ \end{cases},$ (40)
where $\kappa_{1}=\frac{P\sigma_{1}^{-2}\sigma_{2}^{-2}G_{1}G_{2}\bar{\rho}-\sigma_{1}^{-2}G_{1}+\sigma_{2}^{-2}G_{2}}{2\sigma_{1}^{-2}G_{1}}$, and $\kappa_{2}=\frac{P\sigma_{1}^{-2}\sigma_{2}^{-2}G_{1}G_{2}\bar{\rho}+\sigma_{1}^{-2}G_{1}-\sigma_{2}^{-2}G_{2}}{2\sigma_{2}^{-2}G_{2}}$. Accordingly, the NF sum-rate capacity of the BC is given by
$\mathsf{C}_{\mathsf{BC}}=\begin{cases}\log_{2}\big{(}1+P\sigma_{1}^{-2}G_{1}\big{)}&\kappa_{1}\leq 0\\ \log_{2}\big{(}1+P\sigma_{2}^{-2}G_{2}\big{)}&\kappa_{2}\leq 0\\ \log_{2}\big{(}1+\frac{\kappa_{1}+\kappa_{2}+\kappa_{1}\kappa_{2}}{\bar{\rho}}\big{)}&\text{else}\\ \end{cases}.$ (41)
@@
The results can be obtained based on the Karush–Kuhn–Tucker (KKT) conditions of problem (IV-A). ∎
@@
Notably, if the noise power at each UTs is same, for the UPA case, we can deduce that $\lim_{M_{x},M_{z}\rightarrow\infty}\left(p_{1}^{\star},p_{2}^{\star}\right)=\left(\frac{P}{2},\frac{P}{2}\right)$, which indicates that the optimal power allocation strategy tends to an equal power allocation when the number of array element becomes extremely large. On the other hand, in the ULA scenario, the optimal power allocation always depends on the UT’s positions even when $M\rightarrow\infty$.
IV-B Capacity Region
Given a power allocation scheme $\{p_{k}\}_{k=1}^{2}$ with $p_{1}+p_{2}=P$, we can obtain the capacity region of a dual MAC. The capacity region of the BC is the convex hull of all these dual MAC capacity regions [22], which means that the boundary of capacity region of the BC is formed by the corner points of the dual MAC capacity regions, as illustrated in Fig 2. For a given power allocation scheme $\{p_{k}\}_{k=1}^{2}$, the two corner points of the capacity region of the dual MAC, $(\hat{\mathsf{R}}_{1}^{1\rightarrow 2},\hat{\mathsf{R}}_{2}^{1\rightarrow 2})$ and $(\hat{\mathsf{R}}_{1}^{2\rightarrow 1},\hat{\mathsf{R}}_{2}^{2\rightarrow 1})$, are given as follows. When the dirty-paper encoding order is $2\rightarrow 1$, we have
$\hat{\mathsf{R}}_{1}^{1\rightarrow 2}\!=\log_{2}\!\Big{(}1\!+\frac{p_{1}\sigma_{1}^{-2}G_{1}\!+\!p_{1}p_{2}\sigma_{1}^{-2}\sigma_{2}^{-2}G_{1}G_{2}\bar{\rho}}{1+p_{2}\sigma_{2}^{-2}G_{2}}\Big{)},$ (46a)
$\hat{\mathsf{R}}_{2}^{1\rightarrow 2}=\log_{2}\left(1+p_{2}\sigma_{2}^{-2}G_{2}\right).$ (46b)

### v2

@@ III-D Extension to Cases of $K>2$
[…]
${\mathsf{C}}_{\mathsf{MAC}}=\log_{2}\det({\mathbf{I}}_{M}+{\mathbf{H}}{\mathbf{H}}^{\mathsf{H}})=\log_{2}\det({\mathbf{I}}_{K}+{\mathbf{H}}^{\mathsf{H}}{\mathbf{H}}),$ (43)
[…]
${\mathbf{A}}_{K}=\left[\begin{matrix}1+\gamma_{1}\|\mathbf{h}_{1}\|^{2}&\cdots&{\sqrt{\gamma_{1}\gamma_{K}}}\mathbf{h}_{1}^{\mathsf{H}}\mathbf{h}_{K}\\ \vdots&\ddots&\vdots\\ {\sqrt{\gamma_{1}\gamma_{K}}}\mathbf{h}_{K}^{\mathsf{H}}\mathbf{h}_{1}&\cdots&1+\gamma_{K}\|\mathbf{h}_{K}\|^{2}\\ \end{matrix}\right].$ (44)
Note that the diagonal elements of ${\mathbf{A}}_{K}$ are determined by the channel gains ${\mathsf{g}}_{k}\triangleq\lVert{\mathbf{h}}_{k}\rVert^{2}$. Furthermore, the off-diagonal elements, say the $(i,j)$th element $[{\mathbf{A}}_{K}]_{i,j}$ with $i\neq j$, can be written as follows:
$[{\mathbf{A}}_{K}]_{i,j}={\sqrt{\gamma_{i}\gamma_{j}}}\mathbf{h}_{i}^{\mathsf{H}}\mathbf{h}_{j}={\sqrt{\gamma_{i}\gamma_{j}}}\sqrt{{\mathsf{g}}_{i}{\mathsf{g}}_{j}}\frac{\mathbf{h}_{i}^{\mathsf{H}}\mathbf{h}_{j}}{\lVert{\mathbf{h}}_{i}\rVert\lVert{\mathbf{h}}_{j}\rVert}.$ (45)
This fact suggests that the sum-rate capacity when $K>2$ remains a function of the channel gains and the CCF. Therefore, the NF sum-rate capacity can be analyzed in a similar way as in the two-user case. Particularly, as the BS antenna number approaches infinity, it can be concluded that the NF sum-rate capacity approaches its upper limit $\sum_{k=1}^{K}\log_{2}(1+\gamma_{k}{\mathsf{g}}_{k}^{\mathsf{n}})$, which satisfies
$\lim\nolimits_{M_{x},M_{z}\rightarrow\infty}\mathsf{C}_{\mathsf{MAC}}\approx\sum\nolimits_{k=1}^{K}\log_{2}(1+\xi\gamma_{k}/2).$ (46)
The capacity region of the MAC exhibits a polymatroidal structure in $K$-dimensional space, and each of the $K!$ corner points can be attained by employing SIC decoding in a specific message decoding order. The remaining capacity region is realized through time-sharing among these corner points. Specifically, when considering the SIC decoding order $\varpi(K)\rightarrow\varpi(K-1)\rightarrow\ldots\rightarrow\varpi(1)$ with $\{\varpi(k)\}_{k=1}^{K}={\mathcal{K}}$, the achievable rate tuple $({\mathsf{R}}_{1},\ldots,{\mathsf{R}}_{K})$ at the corner point of the capacity region is given by
$\begin{split}\mathsf{R}_{\varpi\left(k\right)}=&\log_{2}\det\big{(}\mathbf{I}_{M}+\gamma_{\varpi\left(k\right)}\mathbf{h}_{\varpi\left(k\right)}\mathbf{h}_{\varpi\left(k\right)}^{\mathsf{H}}\\ &\!\!\times\Big{(}\mathbf{I}+\sum\nolimits_{k^{\prime}<k}{\gamma_{\varpi\left(k^{\prime}\right)}\mathbf{h}_{\varpi\left(k^{\prime}\right)}\mathbf{h}_{\varpi\left(k^{\prime}\right)}^{\mathsf{H}}}\Big{)}^{\!-1}\Big{)}.\end{split}$ (47)
By following steps similar to those in Appendix -A, (47) can be calculated as follows:
$\mathsf{R}_{\varpi\left(k\right)}=\log_{2}(1+\gamma_{\varpi\left(k\right)}(\|\mathbf{h}_{\varpi\left(k\right)}\|^{2}-\mathbf{a}_{k}^{\mathsf{H}}\mathbf{B}_{k}^{-1}\mathbf{a}_{k})),$ (48)
where $\mathbf{a}_{k}=[\mathbf{h}_{\varpi\left(k\right)}^{\mathsf{H}}\mathbf{h}_{\varpi(k-1)},\ldots,\mathbf{h}_{\varpi(k)}^{\mathsf{H}}\mathbf{h}_{\varpi(1)}]^{\mathsf{H}}$, and $\mathbf{B}_{k}=\left[\begin{matrix}1+{\gamma_{\varpi\left(k-1\right)}}\left\|\mathbf{h}_{\varpi\left(k-1\right)}\right\|^{2}&\cdots&{\sqrt{\gamma_{\varpi\left(k-1\right)}\gamma_{\varpi\left(1\right)}}}\mathbf{h}_{\varpi\left(1\right)}^{\mathsf{H}}\mathbf{h}_{\varpi\left(k-1\right)}\\ \vdots&\ddots&\vdots\\ {\sqrt{\gamma_{\varpi\left(k-1\right)}\gamma_{\varpi\left(1\right)}}}\mathbf{h}_{\varpi\left(k-1\right)}^{\mathsf{H}}\mathbf{h}_{\varpi\left(1\right)}&\cdots&1+{\gamma_{\varpi\left(1\right)}}\left\|\mathbf{h}_{\varpi\left(1\right)}\right\|^{2}\\ \end{matrix}\right]$ Eq. (48) can be also expressed in terms of a function of the channel gains and CCFs. Thus, the results derived in the case of $K=2$ can be trivially extended to the case of $K>2$.
@@ IV Broadcast Channel
${\mathsf{R}}_{\varepsilon(k)}=\log_{2}\left(1\!+\!\frac{{\mathbf{h}}_{\varepsilon(k)}^{\mathsf{H}}{\bm{\Sigma}}_{\varepsilon(k)}{\mathbf{h}}_{\varepsilon(k)}}{\sigma_{\varepsilon(k)}^{2}\!+\!\sum_{k^{\prime}<k}{\mathbf{h}}_{\varepsilon(k)}^{\mathsf{H}}{\bm{\Sigma}}_{\varepsilon(k^{\prime})}{\mathbf{h}}_{\varepsilon(k)}}\right).$ (49)
@@
Deriving a closed-form expression for ${\mathsf{C}}_{\mathsf{BC}}$ under the NF model is a challenging task. As a compromise, we continue to consider the two-user case to gain more insights.
IV-A Near-Field Capacity of the Two-User Case
IV-A1 Sum-Rate Capacity
@@
$\begin{split}{\mathsf{C}}_{\mathsf{BC}}=&\max_{{\bm{\Sigma}}_{k}\succeq{\mathbf{0}},\sum_{k=1}^{2}{\mathsf{tr}}({\bm{\Sigma}}_{k})\leq P}\log_{2}\left(1+\frac{{\mathbf{h}}_{1}^{\mathsf{H}}{\bm{\Sigma}}_{1}{\mathbf{h}}_{1}}{\sigma_{1}^{2}}\right)\\ &\qquad\qquad+\log_{2}\left(1+\frac{{\mathbf{h}}_{2}^{\mathsf{H}}{\bm{\Sigma}}_{2}{\mathbf{h}}_{2}}{\sigma_{2}^{2}+{\mathbf{h}}_{2}^{\mathsf{H}}{\bm{\Sigma}}_{1}{\mathbf{h}}_{2}}\right).\end{split}$ (51)
@@
According to (51), solving the broadcast capacity is a non-convex problem, where numerically finding the maximum is nontrivial. However, in [31], a duality is shown to exist between the uplink and downlink which establishes that the dirty-paper capacity region for the BC is equal to the capacity region of the dual MAC (described in (7)). Particularly, in the dual MAC, the UTs in the system have a joint power constraint ($\sum_{k=1}^{K}p_{k}\leq P$) instead of individual constraints ($p_{k}\leq P_{k}$) as in the conventional MAC. Consequently, the BC sum-rate capacity in (51) is equivalent to the following dual MAC capacity:
${\mathsf{C}}_{\mathsf{BC}}=\max_{0\leq p_{k},\sum_{k}p_{k}\leq P}\log_{2}\det\Big{(}{\mathbf{I}}_{M}+\sum\nolimits_{k=1}^{K}\tilde{\gamma}_{k}{\mathbf{h}}_{k}{\mathbf{h}}_{k}^{\mathsf{H}}\Big{)}$
$=\max_{0\leq p_{k},\sum_{k}p_{k}\leq P}\log_{2}(1+\tilde{\gamma}_{1}{\mathsf{g}}_{1}+\tilde{\gamma}_{2}{\mathsf{g}}_{2}+\tilde{\gamma}_{1}{\mathsf{g}}_{1}\tilde{\gamma}_{2}{\mathsf{g}}_{2}(1-\rho)).$ (52)
where $\tilde{\gamma}_{k}=\frac{p_{k}}{\sigma_{k}^{2}}$ denotes the SNR. In this case, the capacity can be found through a convex problem that can be optimally solved. A novel expression for the NF sum-rate capacity of BC is derived in the following theorem.
@@
The optimal power allocation for the dual MAC under NFC is given by
$\left(p_{1}^{\star},p_{2}^{\star}\right)=\begin{cases}\left(P,0\right)&\kappa_{1}\leq 0\\ \left(0,P\right)&\kappa_{2}\leq 0\\ \left(\frac{\kappa_{1}}{\sigma_{2}^{-2}\mathsf{g}_{2}^{\mathsf{n}}},\frac{\kappa_{2}}{\sigma_{1}^{-2}\mathsf{g}_{1}^{\mathsf{n}}}\right)&\mathrm{else}\\ \end{cases},$ (53)
where $\kappa_{1}=\frac{P\sigma_{1}^{-2}\sigma_{2}^{-2}\mathsf{g}_{1}^{\mathsf{n}}\mathsf{g}_{2}^{\mathsf{n}}(1-\rho_{\mathsf{n}})-\sigma_{1}^{-2}\mathsf{g}_{1}^{\mathsf{n}}+\sigma_{2}^{-2}\mathsf{g}_{2}^{\mathsf{n}}}{2\sigma_{1}^{-2}\mathsf{g}_{1}^{\mathsf{n}}(1-\rho_{\mathsf{n}})}$, and $\kappa_{2}=\frac{P\sigma_{1}^{-2}\sigma_{2}^{-2}\mathsf{g}_{1}^{\mathsf{n}}\mathsf{g}_{2}^{\mathsf{n}}(1-\rho_{\mathsf{n}})+\sigma_{1}^{-2}\mathsf{g}_{1}^{\mathsf{n}}-\sigma_{2}^{-2}\mathsf{g}_{2}^{\mathsf{n}}}{2\sigma_{2}^{-2}\mathsf{g}_{2}^{\mathsf{n}}(1-\rho_{\mathsf{n}})}$. Accordingly, the NF sum-rate capacity of the BC is given by
$\mathsf{C}_{\mathsf{BC}}=\begin{cases}\log_{2}\big{(}1+P\sigma_{1}^{-2}\mathsf{g}_{1}\big{)}&\kappa_{1}\leq 0\\ \log_{2}\big{(}1+P\sigma_{2}^{-2}\mathsf{g}_{2}\big{)}&\kappa_{2}\leq 0\\ \log_{2}\big{(}1+{\kappa_{1}+\kappa_{2}+\kappa_{1}\kappa_{2}(1-\rho_{\mathsf{n}})}\big{)}&\text{else}\\ \end{cases}.$ (54)
@@
The results can be derived by applying the Karush–Kuhn–Tucker (KKT) conditions to the right-hand side of (IV-A1). ∎
After obtaining the optimized power allocation $\{p_{k}\}_{k=1}^{2}$ for the dual MAC, we can employ the transformation in [31, Eqs. (8)–(10)] to recover the corresponding downlink covariance matrices that achieve the same rates and the sum-rate. Specifically, for the encoding order $2\rightarrow 1$, the optimal covariance matrices for UT 1 and 2 are, respectively, given by
$\mathbf{\Sigma}_{1}^{\star}=\frac{p_{1}^{\star}\mathbf{\Lambda h}_{1}\mathbf{h}_{1}^{\mathsf{H}}\mathbf{\Lambda}}{\mathbf{h}_{1}^{\mathsf{H}}\mathbf{\Lambda h}_{1}},\ \mathbf{\Sigma}_{2}^{\star}=\frac{p_{2}^{\star}\left(1+\mathbf{h}_{2}^{\mathsf{H}}\mathbf{\Sigma}_{1}^{\star}\mathbf{h}_{2}\right)}{\mathsf{g}_{2}^{\mathsf{n}}}\mathbf{h}_{2}\mathbf{h}_{2}^{\mathsf{H}},$ (55)
where $\mathbf{\Lambda}=\mathbf{I}-\frac{p_{2}^{\star}\mathbf{h}_{2}\mathbf{h}_{2}^{\mathsf{H}}}{1+p_{2}^{\star}\mathsf{g}_{2}^{\mathsf{n}}}$.
We then study the asymptotic NF BC capacity for $M_{x},M_{z}\rightarrow\infty$
@@
Notably, if the noise power at each UTs is same, for the UPA case, we can deduce that $\lim_{M_{x},M_{z}\rightarrow\infty}\left(p_{1}^{\star},p_{2}^{\star}\right)=\left(\frac{P}{2},\frac{P}{2}\right)$, which indicates that the optimal power allocation of the dual MAC tends to an equal power allocation when the number of array element becomes extremely large. On the other hand, in the ULA scenario, the optimal power allocation always depends on the UT’s positions even when $M\rightarrow\infty$.
$\mathsf{C}_{\mathsf{BC}}\simeq\begin{cases}\log_{2}\!\left(\frac{MPA}{4\pi}\max_{k=1,2}\frac{\Psi_{k}}{r_{k}^{2}\sigma_{k}^{2}}\right)\triangleq\mathsf{F}_{\mathsf{BC}}^{\mathsf{s}}&\left(\theta_{1},\phi_{1}\right)=\left(\theta_{2},\phi_{2}\right)\\ \log_{2}\left(\frac{\left(MPA\left(4\pi\right)^{-1}+\Psi_{1}^{-1}r_{1}^{2}\sigma_{1}^{2}+\Psi_{2}^{-1}r_{2}^{2}\sigma_{2}^{2}\right)^{2}}{4\Psi_{1}^{-1}\Psi_{2}^{-1}r_{1}^{2}r_{2}^{2}\sigma_{1}^{2}\sigma_{2}^{2}}-1\right)\triangleq\mathsf{F}_{\mathsf{BC}}^{\mathsf{d}}&\left(\theta_{1},\phi_{1}\right)\neq\left(\theta_{2},\phi_{2}\right)\\ \end{cases}$ (63)
IV-A2 Capacity Region
Given a power allocation scheme $\{p_{k}\}_{k=1}^{2}$ with $p_{1}+p_{2}=P$, we can obtain the capacity region of a dual MAC. The capacity region of the BC is the convex hull of all these dual MAC capacity regions [31], which means that the boundary of capacity region of the BC is formed by the corner points of the dual MAC capacity regions, as illustrated in Fig 4. For a given power allocation scheme $\{p_{k}\}_{k=1}^{2}$, the two corner points of the NF capacity region of the dual MAC, $(\hat{\mathsf{R}}_{1}^{1\rightarrow 2},\hat{\mathsf{R}}_{2}^{1\rightarrow 2})$ and $(\hat{\mathsf{R}}_{1}^{2\rightarrow 1},\hat{\mathsf{R}}_{2}^{2\rightarrow 1})$, are given as follows. When the dirty-paper encoding order is $2\rightarrow 1$, we have
$\hat{\mathsf{R}}_{1}^{2\rightarrow 1}=\log_{2}(1+\tilde{\gamma}_{1}\mathsf{g}_{1}^{\mathsf{n}}),$ (60a)
$\hat{\mathsf{R}}_{2}^{2\rightarrow 1}=\log_{2}\Big{(}1+\frac{\tilde{\gamma}_{2}\mathsf{g}_{2}^{\mathsf{n}}+\tilde{\gamma}_{1}\mathsf{g}_{1}^{\mathsf{n}}\tilde{\gamma}_{2}\mathsf{g}_{2}^{\mathsf{n}}(1-\rho_{\mathsf{n}})}{1+\tilde{\gamma}_{1}\mathsf{g}_{1}^{\mathsf{n}}}\Big{)}.$ (60b)